hello, I need a little help with this problem in the second part. I have tried working the example and I was able to do it, but not my actual problem.

You estimate that by the time you retire in 35 years, you will have accumulated savings of $2.5 million.

a. If the interest rate is 10.5% and you live 15 years after retirement, what annual level of expenditure will those savings support? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Annual expenditure $ b. Unfortunately, inflation will eat into the value of your retirement income. Assume a 6% inflation rate and work out a spending program for your $2.5 million in retirement savings that will allow you to increase your expenditure in line with inflation. What will be your expenditure amount in real terms for each year of your retirement
Answer #1

1.
=2500000*10.5%/(1-1/1.105^15)=338120.0038

2.
real rate=1.105/1.06-1=4.245%
=2500000/1.06^35*4.245%/(1-1/(1+4.245%)^15)=29757.89344